# Two Heads are Better than One: Distilling Large Language Model Features Into Small Models with Feature Decomposition and Mixture

### v2

@@ Introduction
Market making is a core task in financial area, which could provide liquidity to each financial asset (gueant2013dealing). Today, algorithmic systems handle more than 60% of the trading volume in active markets (othman2012automated). Recent breakthroughs in LLM have demonstrated exceptional potential in financial data analysis (brown2020language; radford2019language; li2025m2iv). We conducted an exploratory experiment that shows Gemini-2.5-Pro (team2023gemini) and Llama-3.1 (grattafiori2024llama) perform well in all indicators, surpassing traditional RL methods, as depicted in Figure 1. However, LLM inference speeds cannot meet the demands of real-time trading, where subsecond latency is essential (vaswani2017attention; kaplan2020scaling). Knowledge distillation (KD) (hinton2015distilling; gou2021knowledge; xu2024survey) is a technique used to mitigate latency issues. However, existing approaches focus on distilling knowledge from large LLMs to small LLMs, these small LLMs still exhibit slower inference speeds compared to traditional lightweight models (jiao2019tinybert). There are no cross-architecture distillation methods.
We then conduct an exploratory experiment that directly distills the LLM feature to a small model using traditional distillation methods, resulting in poor performance. This failure results in the single small model lacking the representational capacity to capture an LLM’s deep and high-dimensional complex features. Therefore, we try to decompose complex LLM features into simpler components and assign them to an ensemble of small learning models to improve the effectiveness of small models in learning LLM features. We believe that layers, tasks, and input data types, such as these three variables, have the potential to decouple complex features. To further validate this, we propose a Normalized Fluorescent Probe to analyze the complex feature representations of LLMs (yu2024interpreting). Our analysis shows that with stronger decoupling conditions of these dimensions, the LLM feature exhibits more apparent separation between clusters. Furthermore, a specialization across the model depth is observed: the shallow layers prioritize the prediction of mid-price, the middle layers focus on the spread, and the deep layers are geared towards the total volume(bouchaud2009markets; kyle1985continuous). Building on these insights, we introduce Cooperative Market Making (CMM), a two-stage framework which contains: (1) Orthogonal Feature Decomposition Distillation (OFDD): We decouple LLM features into three dimensions: layer feature hierarchy, task objectives, and data type/market regime to decompose complex feature spaces into specialized clusters (lecun2015deep; tang2024divide). Each cluster is distilled into dedicated lightweight models (ba2014deep). (2) Hájek Projection-based Mixture-of-Experts (Hájek-MoE): We design a projection mechanism to quantify the contribution of each lightweight expert model. (hajek1968asymptotic)
@@
Market making involves the continuous submission of buy and sell orders in the limit order book (gould2013limit) to maximize returns while managing risk (vicente2023automated; amihud1980dealership). Traditional MM frameworks, including those introduced by (avellaneda2008high; gueant2017optimal), are based on mathematical models that often assume static market conditions. In recent years, RL has gained traction as a flexible and adaptive approach for designing MM strategies that can respond to real-time market fluctuations (chan2001electronic; cartea2015algorithmic; patel2018optimizing; zhong2021data; fang2021universal; niu2022metatrader; sun2023mastering), but a significant portion of this research has focused on the refinement of strategies for a single price level (sadighian2019deep). Many RL-based approaches use the mid-price as a dynamic pricing reference, while this reliance on a volatile signal can lead to excessive order cancellations (kumar2020deep).
@@
Recent LLMs, such as PaLM 540b (chowdhery2023palm; li2025cama), pose significant challenges to both inference and fine-tuning due to their substantial computational demands. These dependencies highlight the importance of knowledge distillation. Furthermore, the Chain-of-Thought (wei2022chain) framework has enabled the generation of rich reasoning outputs from teacher models (ho2022large), allowing student models to learn not only the answers but also the reasoning processes. This approach enhances the student model through multitask learning (chu2023multi). In addition, efforts have been made to generate various rationales to improve the consistency of predictions (chen2024towards). Although these systems are designed to leverage knowledge diversity, they remain underexplored.
@@
Mixture of Experts (MoE) was introduced by (jordan1994hierarchical), with distinct experts handling different input sections. It was extended by deep MoE and conditional computation (bertsekas2005convergence). (shazeer2017outrageously) integrated MoE with LSTMs (hochreiter1997long), and Switch Transformers (fedus2022switch) combined MoE with transformers, enabling MoE in architectures like MoE-LLaVA (lin2024moe). The core principle of MoE is to expand the capacity of the model efficiently (dou2019dynamic). Unlike previous MoE techniques (cai2024survey; vats2024evolution) where specialization emerges during training, MoAI (lee2024moai) explicitly defines experts for specific input segments. Recent advances enable training trillion-parameter models in natural language processing (li2021semantics) and computer vision (ochs2015iteratively).
@@
Previous work has shown that the different layers of LLM are characterized at different levels, from shallow to deep (yu2024interpreting). Our normalized fluorescent probe also identified the dividing lines between different layers. Based on this intuition, we make different models learn different parts of the feature during Distillation, which can effectively reduce the difficulty of feature learning. In this way, we initially decompose the features based on the layer. The last three levels that we have discovered with the probe are the output of the separate controls, mid-price, spread, and total volume.
@@
Drawing inspiration from the Hájek projection (hajek1968asymptotic; wager2018estimation), which utilizes a kernel function to map the different spaces into the same space, to manage and leverage various small models effectively, we employ a framework based on the mixture of experts (Hájek-MoE) based on the Hájek projection. Our approach similarly maps the feature spaces of different models into a unified representation space through a kernel function that accepts each model’s features and prediction and outputs each model vector’s coordinates in the mapped kernel space.
@@
The data we use are historical orders and trades on the Shanghai Futures Exchange from July 2021 to July 2022, covering 186 trading days follows IMM (niu2023imm). We select four contracts as 4 datasets: FU, RB, CU, and AG and reconstruct their historical 5-depth limit order books with a 500-millisecond real-time financial period. Our model takes the order book of the previous timestep as input. It directly predicts the mid-price, spread, and volume for the subsequent timestep, which are then used to construct the new order book via a traditional trapezoidal algorithm. We follow the IMM (niu2023imm) to do the dataset split to train and test. We use LLaMA3.1 as our LLM. We use MLP with 2 layers as hajak kernel function. All experiments are conducted on 64 NVIDIA H100 GPUs.
@@
We evaluate the effect of our method on multiple rule-based KD approaches, including FOIC (gavsperov2021market), LIIC (gavsperov2021market), LIIIC (niu2023imm). We also compare with various RL-based MM methods including $RL_{DS}$ (beysolow2019market), $DRL_{OS}$ (chung2022market), IMM (niu2023imm) and CMM. The following performance and risk metrics are employed: 1.Episodic PnL is a natural choice to evaluate the profitability of a MM agent (sutton2018reinforcement). 2.Mean Absolute Position (MAP) accounts for the inventory risk (garleanu2013dynamic). 3.Return Per Trade (RPT) evaluates the agent’s capability of capturing the spread (easley2011microstructure). It is normalized across different markets by the average market spread. 4. PnLMAP defined as PnL divided by the mean absolute position (MAP) in this period (hull2016options). It means the PnL in per unit of inventory and can measure the ability of the agent to profit against inventory risk.

### v3

@@ Introduction
Market making is a core task in financial area, which could provide liquidity to each financial asset (Guéant et al. 2013). Today, algorithmic systems handle more than 60% of the trading volume in active markets (Othman 2012). Recent breakthroughs in LLM have demonstrated exceptional potential in financial data analysis (Brown et al. 2020; Radford et al. 2019; Li et al. 2025a). We conducted an exploratory experiment that shows Gemini-2.5-Pro (Team et al. 2023) and Llama-3.1 (Grattafiori et al. 2024) perform well in all indicators, surpassing traditional RL methods, as depicted in Figure 1. However, LLM inference speeds cannot meet the demands of real-time trading, where subsecond latency is essential (Vaswani 2017; Kaplan et al. 2020). Knowledge distillation (KD) (Hinton 2015; Gou et al. 2021; Xu et al. 2024) is a technique used to mitigate latency issues. However, existing approaches focus on distilling knowledge from large LLMs to small LLMs, these small LLMs still exhibit slower inference speeds compared to traditional lightweight models (Jiao et al. 2019). There are no cross-architecture distillation methods.
We then conduct an exploratory experiment that directly distills the LLM feature to a small model using traditional distillation methods, resulting in poor performance. This failure results in the single small model lacking the representational capacity to capture an LLM’s deep and high-dimensional complex features. Therefore, we try to decompose complex LLM features into simpler components and assign them to an ensemble of small learning models to improve the effectiveness of small models in learning LLM features. We believe that layers, tasks, and input data types, such as these three variables, have the potential to decouple complex features. To further validate this, we propose a Normalized Fluorescent Probe to analyze the complex feature representations of LLMs (Yu and Ananiadou 2024). Our analysis shows that with stronger decoupling conditions of these dimensions, the LLM feature exhibits more apparent separation between clusters. Furthermore, a specialization across the model depth is observed: the shallow layers prioritize the prediction of mid-price, the middle layers focus on the spread, and the deep layers are geared towards the total volume(Bouchaud et al. 2009; Kyle 1985). Building on these insights, we introduce Cooperative Market Making (CMM), a two-stage framework which contains: (1) Orthogonal Feature Decomposition Distillation (OFDD): We decouple LLM features into three dimensions: layer feature hierarchy, task objectives, and data type/market regime to decompose complex feature spaces into specialized clusters (LeCun et al. 2015; Tang et al. 2025). Each cluster is distilled into dedicated lightweight models (Ba and Caruana 2014). (2) Hájek Projection-based Mixture-of-Experts (Hájek-MoE): We design a projection mechanism to quantify the contribution of each lightweight expert model. (Hájek 1968)
@@
Market making involves the continuous submission of buy and sell orders in the limit order book (Gould et al. 2013) to maximize returns while managing risk (Vicente et al. 2023; Amihud and Mendelson 1980). Traditional MM frameworks, including those introduced by (Avellaneda and Stoikov 2008; Guéant 2017), are based on mathematical models that often assume static market conditions. In recent years, RL has gained traction as a flexible and adaptive approach for designing MM strategies that can respond to real-time market fluctuations (Zhong et al. 2021; Fang et al. 2021; Lei et al. 2025; Sun et al. 2023; Chen et al. 2025; Wu et al. 2024; Zhao et al. 2025), but a significant portion of this research has focused on the refinement of strategies for a single price level (Sadighian 2019). Many RL-based approaches use the mid-price as a dynamic pricing reference, while this reliance on a volatile signal can lead to excessive order cancellations (Kumar 2020).
@@
Recent LLMs, such as PaLM 540b (Chowdhery et al. 2023; Li et al. 2025b), pose significant challenges to both inference and fine-tuning due to their substantial computational demands. These dependencies highlight the importance of knowledge distillation. Furthermore, the Chain-of-Thought (Wei et al. 2022) framework has enabled the generation of rich reasoning outputs from teacher models (Ho et al. 2022), allowing student models to learn not only the answers but also the reasoning processes. This approach enhances the student model through multitask learning (Chu et al. 2023). In addition, efforts have been made to generate various rationales to improve the consistency of predictions (Chen et al. 2024). Although these systems are designed to leverage knowledge diversity, they remain underexplored.
@@
Mixture of Experts (MoE) was introduced by (Jordan and Jacobs 1994), with distinct experts handling different input sections. It was extended by deep MoE and conditional computation (Bertsekas et al. 2005). (Shazeer et al. 2017) integrated MoE with LSTMs (Hochreiter 1997), and Switch Transformers (Fedus et al. 2022) combined MoE with transformers, enabling MoE in architectures like MoE-LLaVA (Lin et al. 2024). The core principle of MoE is to expand the capacity of the model efficiently (Dou et al. 2019). Unlike previous MoE techniques (Cai et al. 2024; Vats et al. 2024) where specialization emerges during training, MoAI (Lee et al. 2024) explicitly defines experts for specific input segments. Recent advances enable training trillion-parameter models in natural language processing (Li et al. 2021) and computer vision (Ochs et al. 2015).
@@
Previous work has shown that the different layers of LLM are characterized at different levels, from shallow to deep (Yu and Ananiadou 2024). Our normalized fluorescent probe also identified the dividing lines between different layers. Based on this intuition, we make different models learn different parts of the feature during Distillation, which can effectively reduce the difficulty of feature learning. In this way, we initially decompose the features based on the layer. The last three levels that we have discovered with the probe are the output of the separate controls, mid-price, spread, and total volume.
@@
Drawing inspiration from the Hájek projection (Hájek 1968; Wager and Athey 2018), which utilizes a kernel function to map the different spaces into the same space, to manage and leverage various small models effectively, we employ a framework based on the mixture of experts (Hájek-MoE) based on the Hájek projection. Our approach similarly maps the feature spaces of different models into a unified representation space through a kernel function that accepts each model’s features and prediction and outputs each model vector’s coordinates in the mapped kernel space.
@@
The data we use are historical orders and trades on the Shanghai Futures Exchange from July 2021 to July 2022, covering 186 trading days follows IMM (Niu et al. 2023). We select four contracts as 4 datasets: FU, RB, CU, and AG and reconstruct their historical 5-depth limit order books with a 500-millisecond real-time financial period. Our model takes the order book of the previous timestep as input. It directly predicts the mid-price, spread, and volume for the subsequent timestep, which are then used to construct the new order book via a traditional trapezoidal algorithm. We follow the IMM (Niu et al. 2023) to do the dataset split to train and test. We use LLaMA3.1 as our LLM. We use MLP with 2 layers as hajak kernel function. All experiments are conducted on 64 NVIDIA H100 GPUs.
@@
We evaluate the effect of our method on multiple rule-based KD approaches, including FOIC (Gašperov and Kostanjčar 2021), LIIC (Gašperov and Kostanjčar 2021), LIIIC (Niu et al. 2023). We also compare with various RL-based MM methods including $RL_{DS}$ (Beysolow II and Beysolow II 2019), $DRL_{OS}$ (Chung et al. 2022), IMM (Niu et al. 2023) and CMM. The following performance and risk metrics are employed: 1.Episodic PnL is a natural choice to evaluate the profitability of a MM agent (Sutton 2018). 2.Mean Absolute Position (MAP) accounts for the inventory risk (Gârleanu and Pedersen 2013). 3.Return Per Trade (RPT) evaluates the agent’s capability of capturing the spread (Easley et al. 2011). It is normalized across different markets by the average market spread. 4. PnLMAP defined as PnL divided by the mean absolute position (MAP) in this period (Hull and Basu 2016). It means the PnL in per unit of inventory and can measure the ability of the agent to profit against inventory risk.